OpenStudy (anonymous):

Simplify. n^6 · n^5 ÷ n^4 · n^3 ÷ n^2 · n

hartnn (hartnn):

whatever is multiplied, keep those terms in NUMERATOR, \(\large n^6n^5n^3n\) whatever is divided, keep those terms in DENOMINATOR, \(\large n^4n^2\) now simplify them separately.

OpenStudy (anonymous):

so i take n^6n^5n^3 and times them together

hartnn (hartnn):

yeah, use this: \(\large x^a.x^b=x^{a+b}\) so, what is \(\large n^6n^5n^3n=....?\)

OpenStudy (anonymous):

i got n^90 but i dont think thats right

hartnn (hartnn):

you don't multiply the exponents, you ADD the exponents... try again?

OpenStudy (anonymous):

n^15

hartnn (hartnn):

YES ! :)

hartnn (hartnn):

now, denominator ?

OpenStudy (anonymous):

is it 4+2 or 4*2

hartnn (hartnn):

when you multiply the bases, the exponents always gets ADDED, so it'll be 4+2 :)

OpenStudy (anonymous):

so n^6

hartnn (hartnn):

yup, now you have n^15/n^6 right ? know whqat to do next ?

OpenStudy (anonymous):

i almost wanna say add the exponents or is it subtract

hartnn (hartnn):

when bases get multiply , you ADD the exponents, but when bases get DIVIDED, you \(\color{green}{SUBTRACT}\) the exponents, so subtract :) \(\large \dfrac{x^a}{x^b}=x^{a-b}\)

OpenStudy (anonymous):

so the answer is n^9

hartnn (hartnn):

you are \(\huge \checkmark\) :)

OpenStudy (anonymous):

XD thank you for your help

hartnn (hartnn):

\[ \begin{array}l\color{red}{\text{w}}\color{orange}{\text{e}}\color{#e6e600}{\text{l}}\color{green}{\text{c}}\color{blue}{\text{o}}\color{purple}{\text{m}}\color{purple}{\text{e}}\color{red}{\text{ }}\color{orange}{\text{^}}\color{#e6e600}{\text{_}}\color{green}{\text{^}}\color{blue}{\text{}}\end{array} \]